OpenStudy (anonymous):

solve 2x - 5y = -11 & solve 3x + 13y = 4

OpenStudy (anonymous):

find x and y?

OpenStudy (anonymous):

OH! alright, I suggest you use either 'substitution' or 'elimination'. In substitution' case, you need to find a simple value for y, using basic arithmetic and the first equation in the system. For example: 2x-5y+11=0, divide all of the values by 2 to have a simple value for x. [x-2.5y+5.5=0, or x=2.5y-5.5]. Now, you can plug in your value for x as x into the other equation, or: 3(2.5y-5.5)+13y=4, and solve for y. Once you know the value of y, you can plug it into the first equation, or: 2x-5(y{when you know it, at least. ;)}) = -11 And solve to have your value of x!